# An Iterative Task-Driven Framework for Resilient LiDAR Place Recognition in Adverse Weather

### v1

@@ Abstract
LiDAR place recognition (LPR) plays a vital role in autonomous navigation. However, existing LPR methods struggle to maintain robustness under adverse weather conditions such as rain, snow, and fog, where weather-induced noise and point cloud degradation impair LiDAR reliability and perception accuracy. To tackle these challenges, we propose an Iterative Task-Driven Framework (ITDNet), which integrates a LiDAR Data Restoration (LDR) module and a LiDAR Place Recognition (LPR) module through an iterative learning strategy. These modules are jointly trained end-to-end, with alternating optimization to enhance performance. The core rationale of ITDNet is to leverage the LDR module to recover the corrupted point clouds while preserving structural consistency with clean data, thereby improving LPR accuracy in adverse weather. Simultaneously, the LPR task provides feature pseudo-labels to guide the LDR module’s training, aligning it more effectively with the LPR task. To achieve this, we first design a task-driven LPR loss and a reconstruction loss to jointly supervise the optimization of the LDR module. Furthermore, for the LDR module, we propose a Dual-Domain Mixer (DDM) block for frequency-spatial feature fusion and a Semantic-Aware Generator (SAG) block for semantic-guided restoration. In addition, for the LPR module, we introduce a Multi-Frequency Transformer (MFT) block and a Wavelet Pyramid NetVLAD (WPN) block to aggregate multi-scale, robust global descriptors. Finally, extensive experiments on the Weather-KITTI, Boreas, and our proposed Weather-Apollo datasets demonstrate that, demonstrate that ITDNet outperforms existing LPR methods, achieving state-of-the-art performance in adverse weather. The datasets and code will be made publicly available at https://github.com/Grandzxw/ITDNet.
@@ I Introduction
LiDAR Place Recognition (LPR) is a fundamental prerequisite for autonomous navigation, enabling robots to accurately localize within a global map and navigate effectively [1, 2, 3]. While most LPR systems exhibit robust performance under ”clear” weather conditions, their reliability deteriorates significantly in adverse weather scenarios, which are common in real-world applications [4, 5]. Conditions such as snow, rain, and fog introduce varying levels of weather-induced noise and cause partial point cloud degradation in LiDAR scans, reducing the discriminability of point cloud features and impairing the overall performance of LiDAR perception tasks [6, 7, 8]. Therefore, effectively preprocessing LiDAR data while restoring LiDAR data structural consistency is essential for achieving accurate place recognition under adverse weather conditions, which has become a critical challenge for autonomous robotic navigation systems.
@@
In recent years, researchers have developed various preprocessing techniques to enhance LiDAR perception performance in challenging environments [9, 10, 11]. Among these, point cloud denoising is the most prevalent approach, as it effectively removes noise points, mitigating their impact on downstream tasks such as semantic segmentation and object detection [11, 12]. However, existing point cloud denoising methods focus solely on noise removal without restoring the structural consistency of LiDAR data. This limitation is particularly problematic for the LPR task, which relies on the similarity of global descriptors between query and database frames for accurate recognition. Meanwhile, although LiDAR data preprocessing and LPR are both essential for robust localization, they are typically treated as independent tasks [13, 14, 5]. In most cases, as shown in Fig. 1(b), preprocessing methods are first applied to refine point cloud quality, and the processed data is subsequently fed into a pre-trained LPR model for inference and evaluation [5, 14]. However, this separate optimization overlooks the intrinsic relationship between the two tasks. LiDAR preprocessing focuses on low-level fidelity, while LPR targets high-level scene understanding, leading to suboptimal accuracy and higher training cost. As a result, establishing a positive correlation between LiDAR data preprocessing and the LPR task remains an open challenge.
@@
Figure 2: (a) FFT operation: The central frequency spectrum shows the FFT results of degraded and clean point clouds. After swapping the low-frequency components, the IFFT reconstructs the degraded point cloud, reducing noise and enhancing clarity. (b) DWT operation: The restored image is decomposed into four directional sub-bands (LL, LH, HL, HH) via DWT. The LL sub-band captures the global structure, while LH, HL, and HH preserve fine-grained details.
@@
To address the aforementioned challenges, we propose a unified iterative learning framework that alternately trains the LiDAR Place Recognition (LPR) and LiDAR Data Restoration (LDR) tasks, both of which are image-based, in an end-to-end manner across alternating epochs. In our framework, the LDR module progressively generate higher-quality point cloud inputs for the downstream LPR task, while the LPR module provides pseudo-labels based on global feature representations, these pseudo-labels are then used to formulate an LPR task-driven loss, which supervises the training of the LDR module, guiding the LiDAR data restoration process to better align with LPR performance. Additionally, we propose novel LDR and LPR modules that use frequency-domain representations to improve LiDAR understanding, with noise appearing as high-frequency components and scene structure captured in low frequencies. As shown in Fig. 2(a), we apply Fast Fourier Transform (FFT) in the LDR module, where the central region of the frequency spectrum represents low frequencies, and the edges correspond to high frequencies. The degraded point cloud exhibits sparsity in the low-frequency region, whereas the clean point cloud shows the opposite. By swapping the low-frequency components, the Inverse Fast Fourier Transform (IFFT) of the degraded point cloud reduces noise and enhances clarity. Therefore, the FFT operator effectively separates the ground truth from degradation in the frequency domain. For the LPR module, we use wavelet transform, as shown in Fig. 2(b). The wavelet transform effectively isolates the scene structure in the restored point cloud, extracting multi-scale features, which are crucial for place recognition tasks. The main contributions of this work can be summarized as follows:
@@
  • We propose a novel end-to-end optimization framework that adopts an iterative learning strategy to alternately train two tasks: LPR and LDR. To encourage positive interaction between the two tasks during training, we introduce an LPR task-driven loss, enabling the LPR module to supervise the LDR module. In turn, the LDR module improves the quality of LiDAR data, providing better inputs for downstream place recognition.
  • We design novel modules for both the LDR and LPR tasks. The LDR module incorporates a DDM block to effectively fuse frequency and spatial features via FFT, along with an SAG block that injects global contextual cues to enhance restoration accuracy. For the LPR module, we employ a wavelet pyramid to decompose multi-scale feature maps and introduce an MFT block for direction-aware modeling across different frequency sub-bands. Finally, a WPN block aggregates the multi-level representations into a discriminative global descriptor for robust place recognition.
@@
  • Extensive experiments conducted on both the real-world Boreas dataset, Weather-KITTI, and our newly introduced Weather-Apollo datasets demonstrate that the proposed ITDNet model achieves state-of-the-art performance in place recognition under adverse weather conditions, outperforming existing approaches. To facilitate further research, our code has been open-sourced and made publicly available.
@@ II-A LiDAR Place Recognition Methods
LiDAR Place Recognition (LPR) is essential for enhancing long-term navigation accuracy in simultaneous localization and mapping (SLAM) systems. Early methods focused on designing robust hand-crafted global descriptors, categorized into BEV-based [15, 2], discretization-based [16], and point-based approaches [17, 18]. Recently, learning-based global descriptors have become dominant, divided into point-based and image-based methods. PointNetVLAD [19] combines PointNet [20] and NetVLAD [21] for end-to-end global descriptor generation. LPS-Net [22] introduces Bidirectional Perception Units (BPUs) to enhance context aggregation. OverlapTransformer [23] estimates scan overlap via range image matching using a Siamese network, while CVTNet [1] fuses BEV and range image features through cross-attention. Although these methods perform well under clear weather, their robustness significantly degrades in adverse conditions.
@@ II-B LiDAR Data Preprocessing Methods
Adverse weather degrades LiDAR performance by introducing noise and causing partial point cloud corruption. Preprocessing methods mainly include statistical and deep learning approaches [24, 25]. Statistical methods remove noise based on geometric patterns but often fail in complex environments and are computationally intensive for large datasets [26, 27, 28, 29, 30]. Deep learning methods learn noise distributions from data and can be divided into image-based and 3D-based approaches [11, 10, 31]. Image-based methods [32, 10] project 3D point clouds into 2D, risking loss of spatial information, while 3D-based methods [11] directly process raw point clouds, better preserving geometric structures. Although denoising improves downstream tasks like semantic segmentation and detection, its impact on LPR remains limited. Unlike these tasks, LPR demands structural consistency between query and database scans for accurate recognition.
II-C Multi-Task LiDAR SLAM
Multi-task LiDAR SLAM improves SLAM robustness by integrating multiple perception tasks. Dynamic SLAM [33, 34] uses object detection to mitigate the influence of dynamic objects. SGSR-Net [35] introduces structure-aware semantic guidance for indoor SLAM. DLC-SLAM [14] jointly addresses LiDAR denoising and place recognition with two separate networks. However, existing methods often treat tasks independently, overlooking their intrinsic relationships during training and evaluation. To address this, we propose an end-to-end framework that jointly trains LiDAR restoration and LPR, strengthening their task-level synergy.
@@ III Methodology
In this section, we first present the overall architecture of the proposed ITDNet. Then, we provide a detailed description of the LiDAR Data Restoration (LDR) module and the LiDAR Place Recognition (LPR) module. Finally, we introduce the iterative training strategy used for optimizing ITDNet.
@@ III-A The Overall Architecture of ITDNet
Fig. 3(b) and Fig. 3(c) respectively illustrate the detailed network architectures of the LDR and LPR modules. Given a point cloud $\mathbf{P}_{0}\in\mathbb{R}^{N\times 4}$, we project it into a 2D range image $\mathbf{I}_{0}\in\mathbb{R}^{H\times W\times C}$ using spherical projection [23], where $H\times W$ is the spatial resolution and $C$ encodes attributes such as distance $d$ and intensity $i$. In the LDR module, the input $\mathbf{I}_{0}$ is first processed by a convolution layer to extract initial features $\mathbf{F}_{0}\in\mathbb{R}^{H\times W\times C}$. It then passes through a three-stage U-Net encoder-decoder framework, where each layer integrates a DDM block. The DDM block consists of a Frequency Mixer (FMX) layer for capturing global structures using FFT and a Spatial Mixer (SMX) layer for refining local details via depthwise convolutions and pointwise convolutions. The encoder downsamples feature maps, modeling multi-scale receptive fields and producing a compact latent representation $\mathbf{F}_{\text{d}}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$. A DDM block at the bottleneck further enhances spatial correlations. In the decoder, $\mathbf{F}_{\text{d}}$ first passes through an SAG block, which learns noise semantic characteristics. The learned noise is integrated into $\mathbf{F}_{\text{d}}$ and refined through DDM blocks. The decoder progressively upsamples the features, refining spatial details, while skip connections preserve both structure and fine details, ultimately generating the restored feature $\mathbf{F}_{\text{r}}\in\mathbb{R}^{H\times W\times C}$. For the LPR module, the restored feature $\mathbf{F}_{\text{r}}$ is first processed through convolutional layers and a transformer to extract multi-scale spatial features while preserving global context. These features are further refined by the MFT block, which performs direction-aware modeling across different frequency sub-bands, capturing both local structures and long-range dependencies. Finally, the WPN block aggregates the multi-level representations across scales, producing a discriminative global descriptor.
@@ III-B LiDAR Data Restoration module
$[\mathbf{F}{{}_{\text{r}}},\mathbf{F}{{}_{\text{i}}}]=\text{FFT}(\mathbf{F}_{\text{0}})$ (6)
$\mathbf{A}=\sigma(\text{Conv}([\mathbf{F}{{}_{\text{r}}},\mathbf{F}{{}_{\text{i}}}]))$ (7)
$\mathbf{F}_{\text{f}}=\text{IFFT}(\mathbf{A}\odot[\mathbf{F}{{}_{\text{r}}},\mathbf{F}{{}_{\text{i}}}])+\mathbf{F}_{\text{0}}$ (8)
@@
$\mathbf{F}_{\text{s1}}={\delta}(\text{PWConv}({\delta}(\text{DWConv}(\mathbf{F}_{\text{f}}))))+\mathbf{F}_{\text{f}}$ (9)
$\mathbf{F}_{\text{s}}={\delta}(\text{PWConv}({\delta}(\text{DWConv}(\mathbf{F}_{\text{s1}}))))+\mathbf{F}_{\text{s1}}$ (10)
@@
$\mathbf{e}=\text{GAP}(\mathbf{F}_{\text{in}})$ (11)
@@
$\mathbf{F}^{\prime}=\text{Conv}\left(\sum_{s\in\mathcal{S}}\text{Interp}_{s}\left(\sum_{i=1}^{N}\mathbf{W}_{i}^{s}\mathbf{F}_{i}^{s}\right)\right)$ (13)
@@
Loss Function. To better align the LDR module optimization with the objectives of the LPR task, we introduce an LPR-supervised task-driven loss. Specifically, in each iteration, the LPR model trained in the previous epoch extracts global feature descriptors as pseudo-labels. We then enforce consistency between the restored and clean point clouds by minimizing the KL divergence between their corresponding global descriptors. The loss function $\mathcal{L}{{}_{\text{LDR}}}$ for the LDR module is defined as follows:
=$\mathcal{L}{{}_{\text{LDR}}}=\mathcal{L}{{}_{\text{REC}}}+\lambda\mathcal{L}_{\text{LTD}}$ (14)
@@
$=\sum S\left(f_{\text{P}}(\hat{\mathbf{p}}_{j})\right)\log\frac{S\left(f_{\text{P}}(\hat{\mathbf{p}}_{j})\right)}{S\left(f_{\text{P}}(\mathbf{p}_{j})\right)}$ (17)
Where $\mathcal{L}_{\text{REC}}$ denotes the reconstruction loss and $\mathcal{L}_{\text{LTD}}$ represents the LPR task-driven loss. The parameter $\lambda$ acts as a weighting factor to balance their contributions during training. Specifically, $d_{j}$ and $i_{j}$ are the ground-truth distance and intensity values in the clean LiDAR range image, while $\hat{d}_{j}$ and $\hat{i}_{j}$ are the corresponding restored predictions. $N$ is the number of elements. $S$ denotes the softmax operation, and $f_{\text{P}}$ refers to the LPR module trained in the previous epoch. $\hat{\mathbf{p}}_{j}$ and $\mathbf{p}_{j}$ represent the input features of the restored and clean point clouds, respectively. This loss function $\mathcal{L}_{\text{LDR}}$ not only penalizes absolute deviations between the restored and clean measurements but also enforces global feature consistency through the LPR module. The task-driven LPR loss leverages KL divergence to measure the discrepancy between the global descriptors of the restored and clean point clouds, ensuring that the restored features closely align with those of clean data. By minimizing $\mathcal{L}_{\text{LDR}}$, the LDR module is effectively guided to restore degraded LiDAR inputs in a manner that directly benefits place recognition, bridging the gap between independent restoration and recognition training.

### v2

@@ Abstract
LiDAR place recognition (LPR) plays a vital role in autonomous navigation. However, existing LPR methods struggle to maintain robustness under adverse weather conditions such as rain, snow, and fog, where weather-induced noise and point cloud degradation impair LiDAR reliability and perception accuracy. To tackle these challenges, we propose an Iterative Task-Driven Framework (ITDNet), which integrates a LiDAR Data Restoration (LDR) module and a LiDAR Place Recognition (LPR) module through an iterative learning strategy. These modules are jointly trained end-to-end, with alternating optimization to enhance performance. The core rationale of ITDNet is to leverage the LDR module to recover the corrupted point clouds while preserving structural consistency with clean data, thereby improving LPR accuracy in adverse weather. Simultaneously, the LPR task provides feature pseudo-labels to guide the LDR module’s training, aligning it more effectively with the LPR task. To achieve this, we first design a task-driven LPR loss and a reconstruction loss to jointly supervise the optimization of the LDR module. Furthermore, for the LDR module, we propose a Dual-Domain Mixer (DDM) block for frequency-spatial feature fusion and a Semantic-Aware Generator (SAG) block for semantic-guided restoration. In addition, for the LPR module, we introduce a Multi-Frequency Transformer (MFT) block and a Wavelet Pyramid NetVLAD (WPN) block to aggregate multi-scale, robust global descriptors. Finally, extensive experiments on Weather-KITTI, Boreas, and our proposed Weather-Apollo datasets demonstrate that, ITDNet outperforms existing LPR methods, achieving state-of-the-art performance in adverse weather.
@@ I Introduction
LiDAR Place Recognition (LPR) is a fundamental prerequisite for autonomous navigation, enabling robots to accurately localize within a global map and navigate effectively [1, 2, 3, 4, 5]. While most LPR systems exhibit robust performance under ”clean” weather conditions, their reliability deteriorates significantly in adverse weather scenarios, which are common in real-world applications [6, 7]. Conditions such as snow, rain, and fog introduce varying levels of weather-induced noise and cause partial point cloud degradation in LiDAR scans, reducing the discriminability of point cloud features and impairing the overall performance of LiDAR perception tasks [8, 9, 10]. Therefore, effectively preprocessing LiDAR data while restoring it’s structural consistency is essential for achieving accurate place recognition under adverse weather conditions, which has become a critical challenge for autonomous robotic navigation systems.
@@
In recent years, researchers have developed various preprocessing techniques to enhance LiDAR perception performance in challenging environments [11, 12, 13]. Among these, point cloud denoising is the most prevalent approach, as it effectively removes noise points, mitigating their impact on downstream tasks such as semantic segmentation and object detection [13, 14]. However, existing point cloud denoising methods focus solely on noise removal without restoring the structural consistency of LiDAR data. This limitation is particularly problematic for the LPR task, which relies on the similarity of global descriptors between query and database frames for accurate recognition. Meanwhile, although LiDAR data preprocessing and LPR are both essential for robust localization, they are typically treated as independent tasks [15, 7]. In most cases, as shown in Fig. 1(b), preprocessing methods are first applied to refine point cloud quality, and the processed data is subsequently fed into a pre-trained LPR model for inference and evaluation [7, 15]. However, this separate optimization overlooks the intrinsic relationship between the two tasks [16, 17]. LiDAR preprocessing focuses on low-level fidelity, while LPR targets high-level scene understanding, leading to suboptimal accuracy and higher training cost. As a result, establishing a positive correlation between LiDAR data preprocessing and the LPR task remains an open challenge.
@@
Figure 2: (a) FFT operation: The central frequency spectrum shows the FFT results of degraded and clean point clouds. After swapping the low-frequency components, the IFFT reconstructs the degraded point cloud, reducing noise and enhancing clarity. (b) DWT operation: The restored image is decomposed into four directional sub-bands (LL, LH, HL, HH) via DWT. The LL sub-band captures the global scene structure, while the LH, HL, and HH preserve fine-grained details.
@@
To address the aforementioned challenges, we propose a unified iterative learning framework that alternately trains the LPR and LDR tasks, both of which are image-based, in an end-to-end manner across alternating epochs. In our framework, the LDR module progressively generate higher-quality point cloud inputs for the downstream LPR task, while the LPR module provides pseudo-labels based on global feature representations. These pseudo-labels are then used to formulate an LPR task-driven loss, which supervises the training of the LDR module, guiding feature learning to better align the restoration process with LPR performance. Additionally, we propose frequency-domain blocks for the LDR and LPR modules to suppress noise and enhance structural feature modeling, leveraging the property that noise primarily resides in high-frequency bands while scene structures are encoded in low-frequency components. As shown in Fig. 2(a), we apply the Fast Fourier Transform (FFT) in the LDR module, where low-frequency components, highlighted by the red ellipse, are concentrated near the center of the spectrum, while high-frequency components appear toward the edges. The degraded point cloud exhibits weakened energy in the low-frequency region, whereas the clean point cloud shows the opposite. By swapping the low-frequency components, the Inverse FFT (IFFT) applied to the degraded point cloud effectively suppresses noise and enhances clarity. Therefore, the FFT operator effectively isolates degradation artifacts in the frequency domain. For the LPR module, we use wavelet transform, as shown in Fig. 2(b). The wavelet transform effectively isolates the scene structure in the low-frequency LL sub-band of the restored point cloud and extracts multi-scale features, which are crucial for place recognition tasks. The main contributions of this work can be summarized as follows:
@@
  • We propose an end-to-end optimization framework that jointly trains the LPR and LDR modules via an iterative learning strategy. To encourage task-level interaction during training, we introduce an LPR task-driven loss, enabling the LPR module to supervise the LDR module. In return, the LDR module improves LiDAR data quality, providing better inputs for LPR.
  • We design frequency-domain blocks tailored to the LDR module. Specifically, the DDM block performs feature mixing by alternately applying frequency-domain and spatial-domain operations, effectively suppressing high-frequency noise. The SAG block further refines feature restoration through multi-scale semantic guidance.
  • For the LPR module, the MFT block enhances structural representation by applying frequency-guided window attention to multi-scale wavelet sub-bands, capturing both global structures and directional details. The WPN block then aggregates the multi-level features into a compact global descriptor for robust place recognition.
@@
  • Extensive experiments conducted on both the real-world Boreas dataset, Weather-KITTI, and our newly introduced Weather-Apollo datasets demonstrate that the proposed ITDNet model achieves state-of-the-art performance in place recognition under adverse weather conditions. To support further research, we have open-sourced our code.
@@ II-A LiDAR Place Recognition Methods
LiDAR Place Recognition (LPR) is essential for enhancing long-term navigation accuracy in simultaneous localization and mapping (SLAM) systems. Early methods focused on designing robust hand-crafted global descriptors, categorized into BEV-based [18, 2], discretization-based [19], and point-based approaches [20, 21]. Recently, learning-based global descriptors have become dominant, divided into point-based and image-based methods. PointNetVLAD [22] combines PointNet [23] and NetVLAD [24] for end-to-end global descriptor generation. LPS-Net [25] introduces Bidirectional Perception Units (BPUs) to enhance context aggregation. OverlapTransformer [26] estimates scan overlap via range image matching using a Siamese network, while CVTNet [1] fuses BEV and range image features through cross-attention. Although these methods perform well under clean weather, their robustness significantly degrades in adverse conditions.
@@ II-B LiDAR Data Preprocessing Methods
Adverse weather degrades LiDAR performance by introducing noise and causing partial point cloud corruption. Preprocessing methods mainly include statistical and deep learning approaches [27, 28]. Statistical methods remove noise based on geometric patterns but often fail in complex environments and are computationally intensive for large datasets [29, 30, 31, 32, 33]. Deep learning methods learn noise distributions from data and can be divided into image-based and 3D-based approaches [13, 12, 34]. Image-based methods [35, 12] project 3D point clouds into 2D to improve real-time performance, while 3D-based methods [13] directly process raw data to better preserve geometric structure. Although denoising improves downstream tasks such as semantic segmentation and detection, its impact on place recognition remains limited, as the task requires strong structural consistency between query and database scans. To address this, [7] introduces an independent LiDAR restoration network to support place recognition. However, the method treats restoration and recognition as separate tasks without joint optimization, resulting in suboptimal performance.
II-C Task-Coupled Perception in Adverse Weather
Adverse weather conditions degrade sensor data quality and negatively affect downstream perception tasks. Existing perception methods typically incorporate preprocessing modules to enhance input quality and improve perception performance. Based on the degree of task coupling, these methods can be categorized into separate learning and union learning [36, 37]. Separate learning approaches [15, 38, 39, 40] treat preprocessing and perception as independent tasks that are optimized separately. This strategy overlooks their inherent correlation and does not guarantee that the output of preprocessing benefits the downstream perception task, often leading to suboptimal end-to-end performance. In contrast, union learning approaches[16, 17, 41] jointly train and optimize preprocessing and perception within a unified framework. By designing a task-driven optimization mechanism, the relationship between low-level preprocessing and high-level recognition is effectively modeled, guiding preprocessing to better support perception. This enhances task synergy, improves end-to-end performance, and reduces redundancy[42, 43]. While such union learning strategies have primarily been applied to tasks like object detection and semantic segmentation, we are the first to propose an end-to-end framework that jointly optimizes LiDAR restoration and place recognition under adverse weather, strengthening task-level interaction and robustness.
@@ III Methodology
In this section, we first present the overall architecture of the proposed ITDNet. Then, we provide a detailed description of the LiDAR Data Restoration (LDR) module and the LiDAR Place Recognition (LPR) module. Finally, we introduce the iterative training strategy used to optimize the ITDNet.
@@ III-A The Overall Architecture of ITDNet
Fig. 3(b) and Fig. 3(c) respectively illustrate the detailed network architectures of the LDR and LPR modules. Given a point cloud $\mathbf{P}_{0}\in\mathbb{R}^{N\times 4}$, which includes 3D coordinates and intensity, we project it into a 2D range image $\mathbf{I}_{0}\in\mathbb{R}^{H\times W\times C}$ using spherical projection [26]. Here, $H\times W$ is the spatial resolution and $C$ encodes attributes such as distance $d$ and intensity $i$. In the LDR module, the input $\mathbf{I}_{0}$ is first processed by a convolution layer to extract initial features $\mathbf{F}_{0}\in\mathbb{R}^{H\times W\times C}$. It then passes through a three-stage U-Net encoder-decoder framework, where each layer integrates a DDM block. The DDM block consists of a Frequency Mixer (FMX) layer for capturing global structures using FFT and a Spatial Mixer (SMX) layer for refining local details via depthwise convolutions and pointwise convolutions. The encoder downsamples feature maps, modeling multi-scale receptive fields and producing a compact latent representation $\mathbf{F}_{\text{d}}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$. A DDM block at the bottleneck further enhances spatial correlations. In the decoder, $\mathbf{F}_{\text{d}}$ first passes through an SAG block, which learns noise semantic characteristics. The learned noise is integrated into $\mathbf{F}_{\text{d}}$ and refined through DDM blocks. The decoder progressively upsamples the features, refining spatial details, while skip connections preserve both structure and fine details, ultimately generating the restored feature $\mathbf{F}_{\text{r}}\in\mathbb{R}^{H\times W\times C}$. For the LPR module, the restored feature $\mathbf{F}_{\text{r}}$ is first processed through convolutional layers and a transformer to extract multi-scale spatial features while preserving global context. These features are further refined by the MFT block, which performs direction-aware modeling across different frequency sub-bands, capturing both local structures and long-range dependencies. Finally, the WPN block aggregates the multi-level representations across scales, producing a discriminative global descriptor.
@@ III-B LiDAR Data Restoration module
$(\mathbf{F}{{}_{\text{r}}},\mathbf{F}{{}_{\text{i}}})=\text{FFT}(\mathbf{F}_{\text{0}}),$ (6)
$\mathbf{A}=\sigma(\text{Conv}([\mathbf{F}{{}_{\text{r}}},\mathbf{F}{{}_{\text{i}}}])),$ (7)
$\mathbf{F}_{\text{f}}=\text{IFFT}(\mathbf{A}\odot[\mathbf{F}{{}_{\text{r}}},\mathbf{F}{{}_{\text{i}}}])+\mathbf{F}_{\text{0}}.$ (8)
@@
$\mathbf{F}_{\text{s1}}={\delta}(\text{PWConv}({\delta}(\text{DWConv}(\mathbf{F}_{\text{f}}))))+\mathbf{F}_{\text{f}},$ (9)
$\mathbf{F}_{\text{s}}={\delta}(\text{PWConv}({\delta}(\text{DWConv}(\mathbf{F}_{\text{s1}}))))+\mathbf{F}_{\text{s1}}.$ (10)
@@
$\mathbf{e}=\text{GAP}(\mathbf{F}_{\text{in}}),$ (11)
@@
$\mathbf{F}^{\prime}=\text{Conv}\left(\sum_{s\in\mathcal{S}}\text{Interp}_{s}\left(\sum_{i=1}^{N}\mathbf{W}_{i}^{s}\mathbf{F}_{i}^{s}\right)\right).$ (13)
@@
Task-driven Loss Function. To better align the LDR module optimization with the objectives of the LPR task, we introduce an LPR-supervised task-driven loss. Specifically, in each iteration, the LPR model from the previous epoch provides global descriptors as pseudo-labels, and we enforce consistency by minimizing the KL divergence between the descriptors of restored and clean point clouds. The LDR loss $\mathcal{L}{{}_{\text{LDR}}}$ combines a reconstruction loss $\mathcal{L}_{\text{REC}}$ and an LPR task-driven loss $\mathcal{L}_{\text{LTD}}$:
=$\mathcal{L}{{}_{\text{LDR}}}=\mathcal{L}{{}_{\text{REC}}}+\lambda\mathcal{L}_{\text{LTD}},$ (14)
@@
$=\sum S\left(f_{\text{P}}(\hat{\mathbf{p}}_{j})\right)\log\frac{S\left(f_{\text{P}}(\hat{\mathbf{p}}_{j})\right)}{S\left(f_{\text{P}}(\mathbf{p}_{j})\right)}.$ (17)
where the parameter $\lambda$ serves as a weighting factor to balance the contributions of the two loss components during training. In $\mathcal{L}_{\text{REC}}$, $d_{j}$ and $i_{j}$ represent the ground-truth distance and intensity values, while $\hat{d}_{j}$ and $\hat{i}_{j}$ are their corresponding restored predictions. $N$ denotes the total number of pixels. In $\mathcal{L}_{\text{LTD}}$, $S$ denotes the softmax operation, and $f_{\text{P}}$ refers to the LPR module obtained from the previous training epoch. $\hat{\mathbf{p}}_{j}$ and $\mathbf{p}_{j}$ denote the input features extracted from the restored and clean point clouds, respectively, and $D_{\text{KL}}$ is the Kullback–Leibler (KL) divergence, which quantifies the discrepancy between their global descriptors. By minimizing this divergence, $\mathcal{L}_{\text{LTD}}$ encourages the restored global descriptors to closely align with clean data. Ultimately, optimizing $\mathcal{L}_{\text{LDR}}$ guides the LDR module toward generating restorations that effectively support place recognition, bridging the gap between independent restoration and recognition training.